Let f be a function defined as follows:

1 ?:Q−{0}→R, ?(?)=1− .

?

Determine the set ?(?)

?h??? ????h????????? Q ??????? ?={?: ?=?,

1

Write down the set ?(?) by listing the elements as well as in the descriptive form

?∈Z−{0}}
Answer #1

fi Q1407 R such that f(x) = 1- x Now, A = {x:x=n , NE Z - {0} 9. A = {..., 3, 2, 1, 1, 2, 3 ... 9 f(-3) = 1-(-3) = 4 | f(1) =